OpenStudy (anonymous):

Making sure i got this right :) |6+(-14)|+|-3+(-8)| Is -19 Thx!

OpenStudy (anonymous):

so do |6+(-14)| first then do |-3+(-8)| then add the 2 answers

OpenStudy (anonymous):

8+11 =19

OpenStudy (anonymous):

right!!! good job

OpenStudy (anonymous):

finally lol thx its -19 right

OpenStudy (anonymous):

no why is it not -19

jimthompson5910 (jim_thompson5910):

the answer should be positive since you're adding 2 absolute values

jimthompson5910 (jim_thompson5910):

|some number| + |some number| = positive + positive = positive

OpenStudy (anonymous):

Absolute Values are always positive

OpenStudy (anonymous):

ok...unless its outside the abs/ val. bar right?

jimthompson5910 (jim_thompson5910):

yes if you have a negative number outside the abs val bars, then it could be negative

OpenStudy (anonymous):

right so if it was mulitplied by a negative outside or subtracted to create a negative difference

OpenStudy (anonymous):

well i got -19 the first time but than if everything inside the | | is positive than its positive correct?...sorry im trying not to get confused.

jimthompson5910 (jim_thompson5910):

you are correct, examples |-2| = 2 |-33| = 33 |267| = 267 The absolute value of any number will be positive (unless that number is 0)

OpenStudy (anonymous):

but the equation inside the absolute value bars are conducted normally. so for example |3-10|=|-7|=7

OpenStudy (anonymous):

ok guys! thanks for your help...its set in my mind that evrything inside | | is positive!

jimthompson5910 (jim_thompson5910):

no think of || as a function whatever goes in comes out positive

OpenStudy (anonymous):

so take out all the signs inside | | and just do the equation and it will come out positive?

jimthompson5910 (jim_thompson5910):

more like, evaluate everything inside the || first, then make the result positive example |2+4| = |6| = 6 or |2-4| = |-2| = 2 or |30-70| = |-40| = 40

OpenStudy (anonymous):

ok got ya! thx!

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

yw